OpenStudy (anonymous):

doreen uses 13 apples for every two pies she bakes. How many apples will she use for 12 pies? ratio 13/2=?/12

OpenStudy (anonymous):

?=78 bc of 13x6

OpenStudy (anonymous):

your answer would be 78 , set it up as a proportion 13/2 = x/12 13*12 =156 divide by 2 yep gives you 78

OpenStudy (anonymous):

it means she uses 6,5 apples for one pie (13/2) so for 12 pies she will use 12*6,5 which is 78

OpenStudy (anonymous):

We in American use "6.5" rather than "6,5" but answer is quite correc5t.

OpenStudy (anonymous):

okay my apologies :)

OpenStudy (anonymous):

No apology needed! Just thought you would find it of value.